# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v1

@@ Abstract
Recent minimalist modeling has demonstrated that overdamped polar chiral active matter can support an emergent, inviscid Euler turbulence, despite the system’s strictly dissipative microscopic nature. In this letter, we establish the statistical mechanical foundation for the emergent inertial regime by deriving a formal isomorphism between the model’s agent dynamics and the overdamped Langevin equation for disordered Josephson junctions. We identify the trapped agent state as a macroscopic superconducting phase governed by the Adler equation. The validity of this mapping is confirmed analytically by a disorder-broadened Adler-Ohmic crossover in the system’s slip velocity, corresponding to the saddle-node bifurcation of phase-locking systems. These results define the new minimal chiral flocking model as a motile, disordered Josephson array, bridging active turbulence and quantum superconductivity.
@@ I Introduction
However, an important question remains. Our active agent ensemble is demonstrably overdamped and dissipative, yet the model exhibits an inverse, conservative energy cascade. How does a system that should fundamentally decay instead pump energy towards the largest scales? The resolution lies in considering the discrete distribution of agents, without a renormalized fluid element transformation, explicitly as a spatially extended, disordered Josephson junction array. The depinning transition between ’trapped’ and ’running’ states then generates a macroscopic synchronization stiffness, providing an effective inertia.
@@ II Theoretical Foundation
is the classic tilted washboard potential characterized by two competing energy scales, the potential wells $a_{0}R$ (which synchronize agents), and the tilt $\omega_{i}$ (which de-synchronize agents). From Eq. (5) we observe that agents can be ‘trapped’ when $a_{0}R>\omega_{i}$, ‘running’ when $a_{0}R<\omega_{i}$, and ‘marginal’ when $a_{0}R\approx\omega_{i}$. $\eta_{i}(t)$ (as well as the intrinsic $\omega_{i}$) act as sub-gap leakage [11], the background entropy production caused by noise and disorder. These terms drive an irreducible background entropy production [12], ensuring that the emergent inertial fluid retains a finite ’thermal viscosity’ even deep within the phase-locked regime.
@@
where $\Psi_{i}(t)=\Psi(\mathbf{r}_{i}(t),t)$ is the local phase at agent $i$’s location, representing the collective phase evolution, and $\delta\phi_{i}(t)$ represents an agent’s relative phase phase eovlution. Taking the time derivative of Eq. (6) yields,
@@
the dispersion in the agents’ intrinsic frequency, corresponding to inertial fluid’s topography in the shallow water isomorphism derived in Ref. [2]. Consistent this inference, the effective current in Eq. (10) is a driving force that tries to break local synchronization. Setting Eq. (9) to zero and solving for $\delta\phi$ yields the critical current,
@@
which can likewise be derived from considering the potential well depths in the tilted washboard potential (Eq. 5). When $|I_{\text{eff}}|<I_{c}$, the local coupling force $a_{0}R$ is strong enough to trap the agent, which subsequently locks onto a stable fixed angle and swims with the local group. Conversely, when $|I_{\text{eff}}|>I_{c}$, the agent’s intrinsic frustration exceeds the local maximum coupling force. The agent runs or slips relative to the group, with the phase continuously rotating.
Figure 1: A simulated ensemble of polar chiral active matter turbulence. Panels a–d) show end-state snapshots of the entire simulation space, with agent point-cloud (panel a), topological charge (panel b), flow-lines of circulation in order $R^{2}$ (panel c), and agent slip velocity $v_{\text{slip}}$ (panel d), with respective colorscales where applicable. Panel e) shows defect count as a function of normalized simulation time (left $y$-axis) with a $t^{-0.75}$ powerlaw fit shown with a green, dashed line, and the percentage of trapped agents shown in red (right $y$-axis, inverted), using the heuristic definition of a trapped agent having a slip velocity below 0.05 rad/s. Panel f) shows a cross-correlation analysis of the latter, for normalized times >4.3>4.3, using the $t^{-0.75}$ powerlaw to detrend the defect count timeseries. Estimated confidence intervals for the cross-correlation are shown with dashed, black lines. Panels g–j) show inset $v_{\text{slip}}$-snapshots in rapid succession, highlighting ongoing shock merger, agent thermalization, and subsequent re-organization.
In equivalent terms, for Josephson junctions, when $I_{\text{eff}}<I_{c}$ an oscillator famously locks onto a constant value, with $I_{c}$ there being a material quantity dependent on the semiconductor in question. However, whereas the critical current of a Josephson junction dependends on material qualities, $I_{c}=a_{0}R$ is in the present model entirely dynamic and depends on local synchronization. Still, we observe complete isomorphism between Eq. (9) and the Adler equation for a Josephson junction with in-series resistance [15]. When $I_{\text{eff}}>I_{c}$ the Josephson junction cannot hold onto the phase, acquires a voltage, and subsequently produce entropy through dissipative currents. Equivalently, in our model, when an agent runs or slips relative to the group, it acquires the voltage,
@@
$T_{\text{slip}}=\int_{0}^{2\pi};\frac{d\delta\phi}{\dot{\delta\phi}}=\frac{2\pi}{\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}}},$ (14)
where we inserted our Adler equation into the integral (Eq. 9). Consequently, the instantaneous slip velocity is
$v_{\text{slip}}=\frac{2\pi}{T_{\text{slip}}}=\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}}.$ (15)
@@
In equivalent terms, we follow Ref. [15]’s formalism for the ensemble average, $\langle v_{\text{slip}}\rangle$, obtained by integrating Eq. (15) over the probability distribution $P(I_{\text{eff}})$. This accounts for noise as a fluctuating tilt that statistically drives the system across the critical threshold:
$\langle v_{\text{slip}}\rangle=\int_{|I_{\text{eff}}|>I_{c}}P(I_{\text{eff}})\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}},dI_{\text{eff}}.$ (17)
@@
In what follows, we demonstrate that our minimal active matter model effectively realizes a motile Josephson array. By embedding the discrete oscillator dynamics of Ref. [15] into intrinsically motile (self-propelled) agents, we introduce a transport mechanism absent in standard static arrays [7]. In our model, dissipative excitations are physically advected by the running agents themselves. This active transport expels disorder from the ordered bulk and concentrates it into defect cores, segregating the system into a dual-phase fluid. The segregation mimics the thermodynamic state preparation observed in driven optical systems [17], where the interplay of drive and dissipation stabilizes distinct macroscopic phases (see also Ref. [18]). Specifically, we observe (1) an ideal, inviscid Euler fluid of trapped agents with inertial mass proportional to order squared [2], and (2) a dissipative resistive fluid of running agents. This resistive phase acts as the thermodynamic sink that ultimately sustains the cycle, permitting the regeneration of order via a kinetic Turing instability [1].
@@ III Results
Figure 1g–j), showing four prior $v_{slip}$-snapshots of a limited portion of the simulation space, offer an instructional explanation of the model’s core thermodynamic mechanism. An ordered ($R\approx 1$) region shatters into phase slips (active turbulence), generating entropy; this is the shock merger mechanism identified by Ref. [2]. The agents inside this now-disordered patch are then quickly reorganized into the inertial fluid by the kinetic Turing instability [1]. It is this process that acts as a thermodynamic state preparation cycle [17], effectively ’pumping’ the system from the high-entropy running state back into the low-entropy trapped state, sustaining the Onsager dipole.
Figure 1e, f) demonstrate the foregoing dynamic in rigorous terms. Panel e) shows defect count going down with a characteristic scaling of $t^{-0.75}$ [19, 20], identified by Ref. [2] to be caused by shock dynamics. Predictably, the percentage of trapped agents increase from the initial chaos [18] (red line in Figure 1e, right $y$-axis, inverted). However, subtracting the $t^{-0.75}$ scaling from the defect count reveal that the fluctuations are moderately correlated (Figure 1f, cross-correlation). The apparent discrepancy is resolved by recognizing that, as topological defects annihilate, the kinetic energy stored in the vortex cores is released into the bulk (Figure 1g–j). A reduction in topological defects is therefore immediately followed by a transient decrease in the global trapped agent fraction due to noise-induced delocking [21, 15]. This is consistent with a latent heat release during a dynamical phase transition [7], generating a thermal spike that exponentially and momentarily increases the probability of phase slips.
@@
Figure 2: The result of 72 simulations systematically varying the exponent in the $\omega_{i}$ distribution, yielding a systematically varied $\Delta\omega$, the dispersion in intrinsic agent frustration $\omega_{i}$ (inset). The main plot shows observed ensemble average of the slip velocity $\langle v_{\text{slip}}\rangle$ (voltage) against dispersion $\Delta\omega$ (current), in a log-log “I-V” plot. $\langle v_{\text{slip}}\rangle$ is calculated as the mean value of $v_{\text{slip}}$ for all NN agents, collected during four snapshots of the second half of each simulation. A solid black line shows a $4^{\text{th}}$ order Hermite polynomial fit, with a dashed-black line indicating the knee identified by non-linear least squares minimization [22]. Dashed-blue line shows a $2^{\text{nd}}$ order (piece-wise linear) Hermite polynomial fit with slopes (exponent $\beta$) of 0 and 1, with a blue-dotted line indicating the theoretical mean-field critical threshold dispersion, $\Delta\omega_{c}\approx 9\,\text{s}^{-1}$. The observable onset of slip velocity occurs significantly earlier (black-dashed line, at $\Delta\omega\approx 2\,\text{s}^{-1}$), indicating substantial sub-gap leakage current driven by the intrinsic disorder of the array.
@@
We formalize the foregoing description by systematically varying $\langle I_{\text{eff}}\rangle=\Delta\omega$ (Eq. 11) in successive simulations, essentially implementing a Monte-Carlo simulation of the current-voltage (I-V) curve for the motile Josephson array under consideration. Here, we note that the dispersion $\Delta\omega$ is entirely decided by the exponent $n$ of the powerlaw distribution in $\omega_{i}$; $P(\omega)\sim\omega^{-n}$ [1, 2]. We therefore systematically vary $n$ from $0.5$ to $4$. Figure 2 shows the result in a scatterplot of the ensemble-averaged slip velocity $\langle v_{\text{slip}}\rangle$ against $\Delta\omega$. To characterize the macroscopic response function without imposing the rigid constraints of a single-parameter junction model, we fit $\langle v_{\text{slip}}\rangle$ using a shape-preserving $4^{\text{th}}$-order Hermite polynomial [22] (solid black line). This phenomenological approach accommodates the inhomogeneous broadening caused by the structural disorder of the agent array [15].
@@
Crucially, the solid-black fit in Figure 2 reveals a continuous evolution of the local scaling exponent. At very low dispersion ($\Delta\omega\ll\Delta\omega_{c}$), the response is dominated by a noise floor ($\beta=0$), gradually rising towards $\Delta\omega_{c}\approx 9$ s-1. The $4^{\text{th}}$-order Hermite polynomial fit identifies a knee at $\Delta\omega\approx 2$ s-1, and the relatively large spread between these values is consistent with sub-gap leakage currents smoothing out the $\beta=0.5$ transition [15]. As the dispersion exceeds the critical threshold, $\Delta\omega>\Delta\omega_{c}$, the system quickly and strictly adheres to a linear regime, asymptotically recovering the Adler-Ohmic limit ($\beta=1$).
IV Summary & Discussion
In this letter, we have derived a strict isomorphism between a minimalist active matter model to disordered Josephson junctions. In so doing, we have resolved the origin of emergent inertial flows in overdamped chiral flocks [1, 2]. We demonstrate that the system is governed by a tilted washboard potential, where a thermodynamic transition separates resistive ‘running’ agents from the ‘trapped’ population that sustain a self-organizing inertial cascade. This establishes the minimalist active matter model as a distributed, motile Josephson array.
@@ IV Summary & Discussion
The classification of the minimal model as a motile Josephson array is empirically demanded by the system’s macroscopic response function (Figure 2), which traverses the distinct Adler-Ohmic crossover predicted by the resistively shunted junction formalism [16, 7] with soft knee-geometry identified in Ref. [15], the topological fingerprint of the Adler equation’s saddle-node bifurcation, definitively distinguishing the mechanism from standard viscous fluid transitions and justifying the identification of a tilted washboard potential [8]. The subsequent recovery of the linear Ohmic limit ($\beta=1$) at high dispersion ($\Delta\omega$) confirms that Ref. [2]’s discovery of the system’s ’inertial’ phenomenology is rigorously governed by the non-linear phase topology of disordered oscillator arrays.
@@
The formal isomorphism to the standard resistively shunted junction model strictly relies on the polar symmetry of the interaction (flocking) [23]. nematic active matter (half-integer symmetry) would imply a $\pi$-periodic potential [24], mapping to distinct universality classes (e.g., XY2 models [25]) that do not support the integer vortex topology of the Onsager dipole. In contrast, the $2\pi$-periodicity of the polar alignment interaction $\sin(\phi_{i}-\Psi)$ must map directly to the sinusoidal current-phase relation of a superconducting tunnel junction [5, 6].
@@
By bridging chiral active matter turbulence and superconductivity we have identified phase synchronization as a microscopic engine of macroscopic order, establishing the motile Josephson array as a new universality class of driven chiral flocks.

### v2

@@ Abstract
Recent minimalist modeling has demonstrated that overdamped polar chiral active matter can support emergent, inviscid Euler turbulence, despite the system’s strictly dissipative microscopic nature. In this letter, we establish the statistical mechanical foundation for this emergent inertial regime by deriving a formal isomorphism between the model’s agent dynamics and the overdamped Langevin equation for disordered Josephson junctions. We identify the trapped agent state as carrying non-dissipative (phase rigidity) information supercurrents, analogous to a macroscopic superconducting phase governed by the Adler equation. The validity of this mapping is confirmed analytically and empirically by demonstrating a disorder-broadened Adler-Ohmic crossover in the system’s slip velocity, corresponding to the saddle-node bifurcation of phase-locking systems. These results define the new minimal chiral flocking model as a motile, disordered Josephson array, bridging active turbulence and superconductivity.
@@ I Introduction
However, an important question remains. Our active agent ensemble is demonstrably overdamped and dissipative, yet the model exhibits an inverse, conservative energy cascade. How does a system that should fundamentally decay instead pump energy towards the largest scales? The resolution lies in considering the discrete distribution of agents, without a renormalized fluid element transformation, explicitly as a spatially extended, disordered Josephson junction array. The depinning transition between ’running’ and ’trapped’ states then generates a macroscopic synchronization stiffness. The resulting alignment torque between agents allows for an information supercurrent, the non-dissipative transmission of phase gradients across the ensemble.
@@ II Theoretical Foundation
is the classic tilted washboard potential characterized by two competing energy scales, the potential wells $a_{0}R$ (which synchronize agents), and the tilt $\omega_{i}$ (which de-synchronize agents).
In this dynamic, the stochastic noise $\eta_{i}(t)$ and the intrinsic frequency dispersion in $\omega_{i}$ function as sub-gap leakage [11, 12] driving the irreducible background entropy production described by Ref. [13]. This continuous dissipation ensures that the emergent inertial fluid retains a finite ’thermal viscosity’ that prevents the system from freezing into a glass-state, even deep within the phase-locked regime. This mechanism establishes a direct isomorphism to Ref. [14], which demonstrated that the 2D topology of an inherently disordered Josephson array allows for coherent microwave emissions despite significant microscopic inhomogeneity (see also Ref. [15]).
@@
where $\Psi_{i}(t)=\Psi(\mathbf{r}_{i}(t),t)$ is the local phase at agent $i$’s location, representing the collective phase evolution, and $\delta\phi_{i}(t)$ represents an agent’s relative phase phase evolution. Taking the time derivative of Eq. (6) yields,
@@
the dispersion in the agents’ intrinsic frequency, corresponding to inertial fluid’s topography in the shallow water isomorphism derived in Ref. [2]. Consistent with this inference, the effective current in Eq. (10) is a driving force that tries to break local synchronization. Setting Eq. (9) to zero and solving for $\delta\phi$ yields the critical current,
@@
which likewise follows from considering the potential well depths in the tilted washboard potential (Eq. 5). When $|I_{\text{eff}}|<I_{c}$, the local coupling force $a_{0}R$ is strong enough to trap the agent, which subsequently locks onto a stable fixed angle and swims with the local group. Conversely, when $|I_{\text{eff}}|>I_{c}$, the agent’s intrinsic frustration exceeds the local maximum coupling force. The agent runs or slips relative to the group, with the phase continuously rotating.
In equivalent terms, for Josephson junctions, when $I_{\text{eff}}<I_{c}$ an oscillator famously locks onto a constant value. However, whereas the critical current of a Josephson junction depends on material qualities, $I_{c}=a_{0}R$ is in the present model entirely dynamic and depends on local synchronization. Still, we observe complete isomorphism between Eq. (9) and the Adler equation for a Josephson junction with in-series resistance [11]. When $I_{\text{eff}}>I_{c}$ the Josephson junction cannot hold onto the phase, acquires a voltage, and subsequently produces entropy through dissipative currents. Equivalently, in our model, when an agent runs or slips relative to the group, it acquires the voltage,
@@
$T_{\text{slip}}=\int_{0}^{2\pi}\;\frac{d\delta\phi}{\dot{\delta\phi}}=\frac{2\pi}{\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}}},$ (14)
where we inserted the Adler equation into the integral (Eq. 9). Consequently, the instantaneous slip velocity is
$v_{\text{slip}}=\frac{2\pi}{T_{\text{slip}}}=\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}},$ (15)
where we note that whereas Eq. (8) measures an agent’s realized slip velocity, Eq. (15) provides a theoretical prediction for the slip velocity based on the established response function of an ideal Josephson junction; the convergence of the two at high disorder will provide empirical weight behind our central thesis.
@@
In equivalent terms, we follow Ref. [11]’s formalism for single junctions, adapting it for the ensemble average, $\langle v_{\text{slip}}\rangle$, obtained by integrating Eq. (15) over the probability distribution $P(I_{\text{eff}})$. This accounts for noise as a fluctuating tilt that statistically drives the system across the critical threshold:
$\langle v_{\text{slip}}\rangle=\int_{|I_{\text{eff}}|>I_{c}}P(I_{\text{eff}})\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}}\,dI_{\text{eff}}.$ (17)
@@
In what follows, we demonstrate that our minimal active matter model effectively realizes a motile Josephson array. By embedding the discrete oscillator dynamics of Ref. [11] into intrinsically motile (self-propelled) agents, we introduce a transport mechanism absent in standard static arrays [7]. In our model, dissipative excitations are physically advected by the running agents themselves. This active transport expels disorder from the ordered bulk and concentrates it into defect cores, segregating the system into a dual-phase fluid. The segregation mimics the thermodynamic state preparation observed in driven optical systems [15], where the interplay of drive and dissipation stabilizes distinct macroscopic phases (see also Refs. [14, 19]). Specifically, we observe (1) an ideal, inviscid Euler fluid of trapped agents with inertial mass proportional to order squared [2], and (2) a dissipative resistive fluid of running agents. This resistive phase acts as the thermodynamic sink that ultimately sustains the cycle, permitting the regeneration of order via a kinetic Turing instability [1].
Figure 1: A simulated ensemble of polar chiral active matter agents. Panels a–d) show end-state snapshots of the entire simulation space, with agent point-cloud (panel a), topological charge (panel b), flow-lines of circulation in order $R^{2}$ (panel c), and agent slip velocity $v_{\text{slip}}$ (panel d), with respective colorscales where applicable. In panel a), trapped agents are shown in black while running agents are shown in red; corresponding to low and high $v_{\text{slip}}$ respectively, confirmed by panel d). Panel e) shows defect count as a function of normalized simulation time (left $y$-axis) with a $t^{-0.75}$ powerlaw fit shown with a green, dashed line, and the average potential barrier height $\Delta(t)$ (Eq. 18) in red (right $y$-axis, inverted). Panel f) shows a cross-correlation analysis of the latter two timeseries, for normalized times >5>5. Estimated confidence intervals for the cross-correlation are shown with dashed, black lines. Panels g–j) show inset $v_{\text{slip}}$ and trapped/running agent pointclouds, in rapid succession-snapshots, highlighting ongoing shock merger, agent thermalization, and subsequent re-organization. See Video S1 in the Supplementary Materials for a video of this simulation run.
@@ III Results
Figure 1g–j), showing four prior $v_{slip}$-snapshots of a limited portion of the simulation space, offer an instructional explanation of the model’s core thermodynamic mechanism. An ordered ($R\approx 1$) region shatters into phase slips (active turbulence), generating entropy; this is the shock merger mechanism identified by Ref. [2]. The agents inside this now-disordered patch are then quickly reorganized into the inertial fluid by the kinetic Turing instability [1]. It is this process that acts as a thermodynamic state preparation cycle [15], mimicking coherent microwave emissions in disordered Josephson arrays [14], effectively ’pumping’ the system from the high-entropy running state back into the low-entropy trapped state, thereby sustaining the Onsager dipole.
Figure 1e, f) demonstrate the foregoing dynamic in rigorous terms. First, we define the gap excess $\Delta(t)$, a metric derived from the potential barrier height of the overdamped Adler equation [11]. That is, we sum the potential heights for the trapped population,
$\Delta(t)=\sum_{\text{trapped}}\left(a_{0}R-|\omega_{i}-\dot{\Psi}_{i}|\right),$ (18)
where the argument $t$ indicates that the quantity is a timeseries. $\Delta(t)$ thus measures the sum total depths of the local potential wells [8] ($\mathcal{U}_{B}$, termed potential barrier height in Ref. [11]) that confine each trapped agent’s phase.
Panel e) shows topological defect count going down with a characteristic scaling of $t^{-0.75}$ [20, 21], identified by Ref. [2] to be caused by shock dynamics. At the same time, and with a strong negative correlation coefficient of 0.88-0.88 (Figure 1f), the gap excess $\Delta(t)$, or locking strength of the trapped agents [11], increase. The rise in $\Delta(t)$ during shock mergers confirms the topological heat pump mechanism: the annihilation of resistive defects ($\mathcal{U}_{B}\approx 0$) instantaneously throws agents into deep superconducting traps ($\mathcal{U}_{B}\gg 0$), actively lowering the system’s effective phase temperature.
Simultaneously, we observe in Figure 1(a,d) that agents trapped within the Onsager dipole’s vortex cores are continuously phase-slipping. These cores are identified by Ref. [2] as massless ($R\to 0$) sonic black holes. In this limit, the critical current Eq. (12) vanishes, reducing the Adler equation to its trivial kinematic limit ($\dot{\phi}_{i}\approx\omega_{i}$). Consequently, the vortex cores are purely dissipative singularities where the capacity to sustain supercurrents of phase rigidity has collapsed. This decoupling renders the cores unable to transmit phase stiffness or radiate coherent modes in the shallow water formalism, effectively silencing them within the inertial fluid.
@@
Figure 2: The result of 256 simulations systematically varying the exponent in the $\omega_{i}$ distribution, yielding a systematically varied disorder strength $\Delta\omega$ (keeping the peak of the $\omega_{i}$ unchanged at $0.2$ s-1, $\Delta\omega$ is calculated using the standard deviation), shown in the inset panel. The main plot shows observed ensemble average of the slip velocity $\langle v_{\text{slip}}\rangle$ (voltage) against dispersion $\Delta\omega$ (current), in a log-log “I-V” plot. $\langle v_{\text{slip}}\rangle$ is calculated as the mean value of $v_{\text{slip}}$ for all NN agents, collected during four snapshots of the second half of each simulation. A solid black line shows a $4^{\text{th}}$ order Hermite polynomial fit [22], while a dashed-blue line shows a $2^{\text{nd}}$ order (piece-wise linear) Hermite polynomial fit with slopes (exponent $\beta$) of 0 and 1. A vertical blue-dotted line indicates the theoretical mean-field critical threshold dispersion, $\Delta\omega_{c}\approx 6\,\text{s}^{-1}$. The observable onset of slip velocity occurs significantly earlier (around $\Delta\omega\approx 0.5\,\text{s}^{-1}$), indicating substantial sub-gap leakage current driven by the intrinsic disorder of the array.
The I-V Curve
@@
Next, we shall provide definite empirical evidence for our thesis by systematically varying total agent disorder, termed dispersion in Refs. [1, 2]; that is, $\langle I_{\text{eff}}\rangle=\Delta\omega$ (Eq. 11), in successive simulations, essentially implementing a Monte-Carlo simulation of the current-voltage (I-V) curve for our model. Here, we note that $\Delta\omega$ is entirely decided by the exponent $n$ of the powerlaw distribution in $\omega_{i}$; $P(\omega)\sim\omega^{-n}$ [1, 2]. We therefore systematically vary $n$ from $0.5$ to $4$. Figure 2 shows the result in a scatterplot of the ensemble-averaged slip velocity $\langle v_{\text{slip}}\rangle$ against $\Delta\omega$, aggregating 256 simulations. To characterize the macroscopic response function without imposing the rigid constraints of a single-parameter junction model, we fit $\langle v_{\text{slip}}\rangle$ using a shape-preserving $4^{\text{th}}$-order Hermite polynomial [22] (solid black line). This phenomenological approach accommodates the inhomogeneous broadening caused by the structural disorder of the agent array [11].
@@
Crucially, the solid-black fit in Figure 2 reveals a continuous evolution of the local scaling exponent. At very low dispersion ($\Delta\omega\ll\Delta\omega_{c}$), the response is dominated by a noise floor ($\beta=0$), gradually rising towards $\Delta\omega_{c}\approx 6$ s-1. The gradual rise in $\langle v_{\text{slip}}\rangle$ towards 6 s-1 is consistent with sub-gap leakage currents smoothing out the $\beta=0.5$ transition [11]. As the dispersion exceeds the critical threshold, $\Delta\omega>\Delta\omega_{c}$, the system quickly and strictly adheres to a linear regime, asymptotically recovering the Adler-Ohmic limit ($\beta=1$).
IV Discussion
@@
The classification of the minimal model as a disordered Josephson array is empirically demanded by the system’s macroscopic response function (Figure 2), which traverses the distinct Adler-Ohmic crossover predicted by the resistively shunted junction formalism [18, 7] with the disorder-induced rounding identified in Ref. [11], a signature of the Adler equation’s saddle-node bifurcation. This distinguishes the model’s central mechanism from standard viscous fluid transitions and justifies the identification of a tilted washboard potential [8]. The subsequent recovery of the linear Ohmic limit ($\beta=1$) at high-disorder confirms that Ref. [2]’s charcterization of the model’s emergent, inertial superfluid is rigorously predicted by the non-linear phase topology of disordered oscillator arrays.
As the foregoing follows from the alignment term $a_{0}R\sin\delta\phi$ in our model being mathematically indistinguishable from the Josphson energy $E_{J}\sin\psi$, we can interpret the Frank elastic constant in active flocks [23] as phase rigidity that enables information supercurrents. These currents are sustained by agents trapped in a tilted washboard potential maintained by the ensemble, analogous to Cooper pair supercurrents [24].
These findings compel the classification of active chiral matter as a high-temperature superconductor for information. While odd viscosity successfully captures the macroscopic fluidity of the state [25, 26], our results demonstrate that its constitutive rigidity is fundamentally governed by Josephson critical currents $I_{c}$, sustained by the phase stiffness of the interaction network. Consequently, the Josephson inequality $|\omega-\dot{\Psi}|<I_{c}$ imposes a hard physical limit on the flock’s maneuverability: there exists a maximum turning rate beyond which the effective ‘voltage drop’ between agents exceeds the critical threshold, resulting in information loss via dissipative phase slippage.
@@
It is important, however, to stress that the formal isomorphism to the standard resistively shunted junction model strictly relies on the polar symmetry of the interaction (flocking, [27]). nematic active matter (half-integer symmetry) would imply a $\pi$-periodic potential [28], mapping to distinct universality classes (e.g., XY2 models [29]) that do not support the integer vortex topology of the Onsager dipole. In contrast, the $2\pi$-periodicity of the polar alignment interaction, $\sin(\phi_{i}-\Psi)$, must map directly to the sinusoidal current-phase relation of a superconducting tunnel junction [5, 6].
@@
By bridging chiral active matter turbulence and superconductivity we have identified phase synchronization as a microscopic engine capable of creating macroscopic order, elucidating the core thermodynamic mechanism of new universality class for driven chiral flocks [2].
V Conclusion
In this letter, we have derived a strict isomorphism between a minimalist active matter model and resistively shunted junction theory, by recovering the Adler equation in the model’s dynamics (Eq. 9), and subsequently following the formalism developed by Ref. [11] for individual disordered Josephson junctions. We demonstrated that the system is governed by a tilted washboard potential, where a thermodynamic transition separates resistive ‘running’ agents from the ‘trapped’ population, the latter of which explicitly sustain an information supercurrent. As such, we have resolved the origin of emergent inertial flows in overdamped chiral active matter [2]. This establishes the minimalist active matter model as a distributed, disordered, motile Josephson array.